Find the surface area of a cylinder if the height is 9 meters and the radius of a base is 6 meters. Use pie( no way to type sign) equals 3.14 and round to the nearest whole number.
Answer by checkley71(8403)

A BASE=PI*R^2=3.14*6^2=3.14*36=113.04 BASE AREA.
CIRCUMFERANCE=2PI*R=2*3.14*6=37.68
37.68*8=339.12 AREA OF THE CYLINDER SURFACE.
ASSUMING YOU WANT TO INCLUDE THE TOP & BOTTOM THEN:
2*113.04+339.12=565.2 TOTAL SURFACE AREA.